Accelerator Update for January 21, 2009 - January 23, 2009 |
Wednesday Jan 21 The day shift began with Operations monitoring the Tevatron (TeV) store 6746 (117.98E30) and the Recycler stash (223.7E10), and the Antiproton Source stacking (25.7E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the MiniBooNE and NuMI experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams. At 8:19 AM, an Accumulator injection kicker (A:IKIK) failed. Experts investigated. Electron Cooling (ECOOL) experts began a 30-minute trajectory study around 11 AM. TeV experts began a 90-minute end-of-store crystal collimator study at 1:29 PM. Operations terminated store 6746 at 3:07 PM. Operations put the TeV into shot setup at 3:19 PM. At 3:41 PM, Operations reported that a TeV sector C17 separator had sparked. An operator reset the spark counter. At 3:55 PM, two MI kickers (KPS4A and B) faulted. They reset. Operations reported at 3:58 PM that the Recycler had stashed antiprotons seven times during the day shift. Operations established store 6747 at 5:10 PM with an initial luminosity of 333.57E30. At 5:39 PM, Operations reported that the Recycler VSA was not working. It was restarted after turning the program off for 30-minutes. At 6:30 PM, Operations reported a rash of Pbar, MI, P1 transfer line, and NuMI aborts due to some kind of position and RF changes in MI. Operators lowered the beam intensity, retuned, reloaded some autotune programs, and eventually raised the beam intensity back to normal. At 6:54 PM, A:IKIK tripped off. It reset. At 10:44 PM, Operations reported several MI kicker (KPS5S) trips. It reset. Operations reported at 11:51 PM that the Recycler had stashed seven times during the evening shift. Thursday Jan 22 The midnight shift began Operations monitoring store 6747 (131.33E30) and the stash (199.7E10), with Pbar stacking (29.1E10), and with NuMI and MiniBooNE taking beam. At 2:56 AM, a MI kicker (KPS4S) tripped off for the first of four times in four hours. It reset. At 3:19 AM, the P1 transfer line beam permit tripped off due a MI power supply’s current drifting out of its limits. Everything reset. Operations disabled the MiniBooNE beam at 6:32 AM for a scheduled access later in the morning. Operations began preparing MiniBooNE for access at 7:52 AM. Operations reported at 8:01 AM that the Recycler had stashed nine times during the midnight shift. Operations terminated store 6747 at 8:03 AM. Operations put the TeV into shot setup around 8:18 AM. A Recycler expert reported at 8:43 AM that the Pelletron had a software issue. They worked on the problem and reported about an hour later that it was fixed. At 9:20 AM, Operations reported that they were having problems with the Recycler VSA again. MiniBooNE came out of access at 9:47 AM. A Recycler expert began an ECOOL study at 10:03 AM. Operations established store 6748 at 10:14 AM with an initial luminosity of 316.37E30. Operations established beam to MiniBooNE at 10:17 AM. At 12:41 PM, BRF13 and 14 tripped off due to a vacuum pump trip (IPRF23). They reset. Operations turned off the MI beam at 1:59 PM so experts could work on SEXV, KPS5S, and KPS4. Everything was repaired by 3:18 PM. At 3:29 PM, KPS4A and B tripped off on a fire fault for the first of three times in 18 minutes. Experts investigated. Operations reported at 4:01 PM that the Recycler had stashed three times during the day shift. At 4:44 PM, BRF7 and 8 tripped off due to a vacuum pump trip (IPRF17). Everything reset. At 4:51 PM, KPS4A and B tripped off with a no fire indication. Experts investigated. At 6:53 PM, KPS4A and B tripped off when KPS4B’s filament faulted. It reset. At 7:28 PM, KPS4A and B tripped off for the first of eight times in about two hours. They reset. At 9:13 PM, KPD3A tripped off on a filament timing indication. It reset. At 11:03 PM, KPS4A and B tripped off on a fire fault. It reset. Operations reported at 11:57 PM that the Recycler had stashed eight times during the evening shift. Friday Jan 23 The midnight shift began with Operations monitoring store 6748 (80.61E30) and the stash (303.4E10), with Pbar stacking (34.6E10), and with NuMI and MiniBooNE taking beam. At 12:08 AM, BRF14 tripped off. It reset. At 1:45 AM, KPS4A and B tripped off for the first of five times in about two hours. They reset. At 2:45 AM, KPS5S tripped off on a “Thyratron No Fire” indication. It would not reset. Operations contacted an expert. Operations put Pbar into standby mode at 2:45 AM. At 3:29 AM, Operations reported that they could not find or resolve the KPS5S problem. Experts were called in. At 6:50 AM, Operations reported that something had happened out at MI-52 that would not allow plots. There was also some odd behavior with the Recycler. A reboot of the MI-2 front end brought things back, but then when Operations turned beam back on there were many BLM aborts. Operators investigated and discovered a burnt smell coming from a sextupole power supply, which they turned off. A power supply expert was called in. At 7:24 AM, a MI Lambertson power supply (LAM52) tripped off and would not reset. Operations called in an expert. At 7:44 AM, Operations reported a rash of Recycler alarms showing clock errors. Experts reported that TCLK was not stable between MI-30 and MI-40. They reported that they would replace the repeater. Operations reported at 7:56 AM that the Recycler had stashed three times during the midnight shift. The Plans for Friday and the Weekend The plans for today are to fix KPS5S, establish a store, and then resumed stacking, stashing, and storing for today and the rest of the weekend. |
|
More Information Comments and Suggestions |
| last modified 01/23/2009 email Fermilab |
| Security, Privacy, Legal | |